OpenStudy (anonymous):

Which of the following equations represents a perpendicular line to 6x - 2y = 20? y = 3x + 2 y = -3x + 4 y equals one third x minus 3 y equals negative one third x minus 5

OpenStudy (jdoe0001):

what's the slope of => 6x - 2y = 20?

OpenStudy (anonymous):

Do I bring the 6x on the other side with the 20 ?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

isolate "y"

OpenStudy (anonymous):

So it should look like -2y = -6x + 20 ?

OpenStudy (jdoe0001):

yes, now divide each side by -2 so you can solve for "y"

OpenStudy (anonymous):

y = 6/2x + 10 ?

OpenStudy (jdoe0001):

\(\bf -2y = -6x + 20 \implies \cfrac{-2y}{-2}=\cfrac{-6x}{-2} + \cfrac{20}{-2}\)

OpenStudy (anonymous):

So where do I go from there ?

OpenStudy (jdoe0001):

well, what did you get for the slope for that one?

OpenStudy (anonymous):

6/2x I Believe

OpenStudy (jdoe0001):

would you say that 6/2 = 3?

OpenStudy (anonymous):

yes if you divide ! but we are still left with 20/-2

OpenStudy (jdoe0001):

well, woudl you say that 20/-2 = -10?

OpenStudy (anonymous):

yes ! but y = 3x - 10 is not one of my answer choices thats what im trying to figure out .

OpenStudy (jdoe0001):

\(\bf \cfrac{-2y}{-2}\cfrac{-6x}{-2} + \cfrac{20}{-2} \implies y = 3x-10\)

OpenStudy (anonymous):

My answer choices are y = 3x + 2 y = -3x + 4 y equals one third x minus 3 y equals negative one third x minus 5

OpenStudy (jdoe0001):

so the slope for this equation is "3", as you can see from the slope-intercept form any perpendicular line to this line, will have a NEGATIVE RECIPROCAL slope what does that mean? RECIPROCAL of 3 1/3 NEGATIVIZE that -1/3 so any lines with that slope, are PERPENDICULAR to this one

OpenStudy (anonymous):

So which one is the answer ? y = 3x + 2 ?

OpenStudy (jdoe0001):

the line that has that slope in negative reciprocal form :)

OpenStudy (anonymous):

so y equals negative one third x minus 5 is the answer ?

OpenStudy (jdoe0001):

y = -1/3 x -5 ^ slope so, yes, -1/3 is the negative reciprocal of 3

OpenStudy (anonymous):

Can I ask you another question real quick ?

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Is zero consider an even number ?

OpenStudy (jdoe0001):

hmm, I'd say no

OpenStudy (anonymous):

ok Thank You For Your Help .

OpenStudy (jdoe0001):

yw